OpenStudy (anonymous):

find n such that x^(8n + 3)=x^(n+1) is divisible by 23 for any integer x

OpenStudy (anonymous):

\[x ^{8n+3} = x^{n+1}\] the equal sign should be 3 bars not 2

OpenStudy (anonymous):

congruent in what sense? Is it modulus?

OpenStudy (anonymous):

frankly to be honest I haven't got a clue...that's all the question said this is the start of the question I think. the = is congruent and comma start of a new line \[x ^{8n +3} = x^{n+1} (\mod 23), x^{8n+3-(n+1)}=1, x^{7n}=1\]

ganeshie8 (ganeshie8):

u may use \(\equiv\) for congruence..

OpenStudy (anonymous):

so I imagine that x^7n must be congruent with 1 (mod 23)

OpenStudy (anonymous):

\[ x^{8n+3}= x^{7n+2}x^{n+1} \]

OpenStudy (anonymous):

dunno if that gets you anywhere

ganeshie8 (ganeshie8):

\(\mathbb{x ^{8n +3} \equiv x^{n+1} (\mod 23)}\) \(\mathbb{x ^{7n+3} \equiv x (\mod 23)}\) By Fermat, \(\mathbb{7n+3 = 23}\) \(\mathbb{\implies n = 20/7}\) however thats not a integer. fail

ganeshie8 (ganeshie8):

from which chapter is this question from ?

OpenStudy (anonymous):

n doesn't have to be an integer x does I think

ganeshie8 (ganeshie8):

for divisibility+fermat all needs to be integers. not just x

OpenStudy (anonymous):

okay...this is actually my homework so I have no clue...I don't get my lecturer whatsover yet somehow i managed to get most of the homework done.

ganeshie8 (ganeshie8):

this is a number theory problem. may i knw from which chapter/mopule this belongs to ? that info may help in approaching the problem ..

OpenStudy (anonymous):

could it be something in encoding???? and perhaps would euclids help???? but really i am not sure

ganeshie8 (ganeshie8):

Actually you're right \(\large n = \frac{20}{7}\) works perfectly fine here. it doesnt make the exponent of x a fraction cuz 7 cancels out ( 7n)..